# CoSPED: Consistent Soft Prompt Targeted Data Extraction and Defense

### v2

@@ Introduction
Comparison against Prior Works
Discussion
CoSPED VS. Ethicist, CLM
@@ Discussion
Table 5 shows Pythia 1.4B results using tokenizer transferred dataset from Dataset $\mathcal{D}_{2}$ with padding. It is obvious that Pythia series model shows different patterns from GPT-Neo series.
Firstly, increasing soft prompt length in the Pythia model testing continues to improve results. However, it exhibits a similar trend to GPT-Neo, in that soft prompts beyond 100 tokens have a minor effect on improving the Exact Extraction Rate.
Notably, Pythia 1.4B is more resilient to extraction than GPT-Neo 1.3B, achieving lower extraction rates among all similar settings. This could be evidenced from different model structures. Pythia 1.4B is likely more resilient because it uses rotary positional embeddings for better contextual encoding, parallel residual connections for more stable information flow, and a larger intermediate size (8192) enabling richer representations. Unlike GPT-Neo’s alternating global/local attention, Pythia applies full attention consistently, which improves its ability to model prompt-context interactions robustly. These design choices, taken together, make it harder for soft prompts to manipulate Pythia’s outputs effectively.
Table 5: CoSPED Experiment Results on Pythia 1.4B model
@@
A similar trend with GPT-Neo is that increasing prefix length will improve the Exact Extraction Rates in all settings, and also, prefix length has a stronger impact than soft prompt length.
@@
In this section, we discuss methods for preventing data extraction attacks on language models. Recent works (Patil, Hase, and Bansal 2023b) have shown that ROME (Meng et al. 2023a) and MEMIT (Meng et al. 2023b) model editing methods can reduce memory leakage. We adapt ROME for our use case involving prefix-suffix pairs, as privacy extraction attacks exploit memorized training sequences.
@@
ROME as Rank-One Model Editing, adjusts model representations to make specific single-token outputs less likely. However, our scenario uses 50-token prefix inputs with corresponding suffix outputs, requiring adaptation for longer sequences.
We focus on the first suffix token following the prefix. This choice is motivated by: (1) exact-match evaluation requires continuous correct generation from the beginning; (2) autoregressive models exhibit cascade effects where early errors propagate; (3) disrupting the initial token prevents full sequence extraction efficiently. For example, given “Living without an aim is like”, we target the model’s tendency to generate “sailing” from the memorized suffix “sailing without a compass.”
We use delta accuracy to measure output likelihood changes before and after editing. In our defense context, delta accuracy near zero indicates preserved general capabilities while reducing privacy leakage. Negative values demonstrate successful privacy protection.
Model utility is assessed using the LAMBADA dataset
(Paperno et al. 2016), which evaluates long-range language modeling through accuracy and perplexity metrics.
Evaluation Metrics
@@
The LAMBADA benchmark assesses long-range language modeling using two metrics: Accuracy (LB Acc), which measures the proportion of correctly predicted final words in context-rich passages, and Perplexity (LB PPL), which evaluates the model’s fluency and linguistic reasoning. Higher accuracy and lower PPL indicate better language modeling performance.
@@
We evaluate against soft prompt extraction attacks using 1000 test prefix-suffix pairs. Table 7 shows extraction rates:
Table 7: Privacy Leakage Methods Exact Extraction Rates
@@
Our ROME method reduces extraction rates from up to 64% to near zero across both attacks. Defense effectiveness increases with longer token-required matches, successfully disrupting autoregressive generation. The minimal changes in $\delta$ Accuracy and LB evaluations confirm that our modified ROME defense selectively targets memorized content without impacting general capabilities.
@@ Conclusion
In this work, we proposed novel loss functions and decoding strategies that greatly improve extraction performance and enhance understanding of training data leakage risks. Through ablation studies, we identified optimal designs that maximize exact extraction rates, with CoSPED consistently achieving strong results across different metrics.
@@
Although we performed multiple model testing, soft prompt requires retraining. One potential future direction is to do transferability of soft prompt embeddings, referencing in (Kim et al. 2023; Maus et al. 2023). (Maus et al. 2023) proposed a method using Euclidean distance that projects the soft prompt embeddings to the nearest vocabulary, so theoretically it is possible to transfer soft prompt embeddings to hard tokens to be universal.
Moreover, developing more robust defense mechanisms to protect LLMs from soft prompt data extraction attacks is also critical for future work. By analyzing how these attacks exploit vulnerabilities in the model’s architecture and training data, we hope to identify strategies to mitigate or prevent unauthorized data access, making LLM more resilient to such attacks in real-world scenarios.
@@ Appendix A Training Settings
In this section, we provide details of our training settings and parameters for reference. Our hardware environment is Intel(R) Core(TM) i9-11900K 3.50GHz CPU, 3 NVIDIA Corporation GA102 RTX 3090 GPUs, 130GiB of memory, and Ubuntu 22.04.4 LTS.
@@
Smooth loss sets top $N$ at 5. Focal loss in Eq. 1 contains $\alpha$ at 0.6 and $\gamma$ at 2.0. Dynamic loss in Eq. 2b has $N$ at 5, $T_{D}$ at 0.1, and $\alpha$ at 5. Additive loss in Eq.3 has $\alpha$ at 0.7, and Common loss has $\beta$ at 0.2. Self Consistency decoding strategy sets $N_{\text{SCD}}$ at 3, and $\mathcal{D}_{\text{Optimal}}$ at 0.7. The seed is set at 42, and we utilize fp16 for faster training.

### v3

@@ Discussion
Table 4: Experiment Results of Methods and Baselines
@@
Table 5 shows Pythia 1.4B results using tokenized dataset from Dataset $\mathcal{D}_{2}$ with padding. It is notable that Pythia model family shows different patterns from GPT-Neo series.
Firstly, increasing soft prompt length in Pythia model testing continues to improve results. However, it exhibits a similar trend to GPT-Neo, in that soft prompts beyond 100 tokens have a minor effect on improving the Exact Extraction Rate.
Notably, Pythia 1.4B is more resilient to extraction than GPT-Neo 1.3B, achieving lower extraction rates among all similar settings. This could be evidenced because Pythia model uses rotary positional embeddings for better contextual encoding, parallel residual connections for more stable information flow, and a larger intermediate size (8192) enabling richer representations. Unlike GPT-Neo’s alternating global/local attention, Pythia applies full attention consistently, which improves its ability to model prompt-context interactions robustly. These design choices, taken together, make it harder for soft prompts to manipulate Pythia’s outputs effectively.
@@
Table 5: CoSPED Experiment Results on Pythia 1.4B model
Similarly, for both Pythia and GPT-Neo, increasing prefix length improves extract rate across all settings, with prefix length having stronger impact than soft prompt length.
@@
In this section, we discuss methods for preventing data extraction attacks on language models. Recent works (Patil, Hase, and Bansal 2024) have shown that ROME (Meng et al. 2022) and MEMIT [^8] MEMIT: https://memit.baulab.info/ (Meng et al. 2023) model editing methods can reduce memory leakage. We adapt ROME for our use case involving prefix-suffix pairs, as privacy extraction attacks exploit memorized training sequences.
@@
ROME as Rank-One Model Editing, adjusts model representations to make specific single-token outputs less likely. However, our scenario uses 50-token prefix inputs with corresponding 50-token suffix outputs, requiring adaptation for longer sequences.
We focus on the first suffix token following the prefix. This choice is motivated by: (1) exact-match evaluation requires continuous correct generation from initials; (2) autoregressive models exhibit cascade effects where early errors propagate; (3) disrupting initial token strongly prevents full sequence extraction. For example, given “Living without an aim is like”, we target model’s tendency to generate “sailing” from memorized suffix “sailing without a compass.”
Defense Evaluation Metrics
@@
Model utility is assessed using LAMBADA dataset [^9] LAMBADA: https://zenodo.org/record/2630551 (Paperno et al. 2016), which evaluates long-range language modeling through accuracy and perplexity metrics. LAMBADA Benchmark assesses long-range language modeling using two metrics: Accuracy (LB Acc), measuring the proportion of correctly predicted final words in context-rich passages, and Perplexity (LB PPL), evaluating model’s fluency and linguistic reasoning. Higher accuracy and lower PPL indicate better language modeling performance.
Early Stopping Criteria
@@
We evaluate against soft prompt extraction attacks using 1000 test prefix-suffix pairs. Table 7 shows Exact Extraction Rates for CoSPED and Ethicist before and after ROME defense on GPT-Neo 1.3B model.
@@
Table 7: Privacy Leakage Methods Exact Extraction Rates
Our ROME method reduces extraction rates from up to 64% at $\text{ER}_{50}$ to near zero across both attacks. Defense effectiveness increases with longer token-required matches, successfully disrupting autoregressive generation. The minimal changes in $\delta$ Accuracy and LB evaluations confirm that our modified ROME defense selectively targets memorized content without impacting general capabilities.
@@ Conclusion
In this work, we proposed novel loss functions and decoding strategies that improve extraction performance and enhance the understanding of data leakage risks. Through ablation studies, we identified optimal designs that maximize exact extraction rates, with CoSPED consistently achieving strong results across different metrics, proving the consistency between generated outputs and memorized training data.
@@
Although we performed multiple model tests, soft prompt requires retraining. One potential improvement is to explore transferability of soft prompt embeddings, referenced in (Kim et al. 2023; Maus et al. 2023). Specifically, projecting soft embeddings to vocabulary tokens via Euclidean distance (Maus et al. 2023) provides theoretical support for universal usage through hard token conversion.
Moreover, developing more robust defense mechanisms to protect LLMs from soft prompt data extraction attacks is also critical for future work. By analyzing how these attacks exploit vulnerabilities in the model’s architecture and training data, we hope to identify strategies to mitigate or prevent unauthorized data access, making LLMs more resilient to such attacks in real-world scenarios.
@@ Appendix A Training Settings
In this section, we provide details of our training settings and parameters for reference. Our hardware environment is Intel(R) Core(TM) i9-11900K @ 3.50GHz CPU, 3 NVIDIA Corporation GA102 RTX 3090 GPUs, 130GiB of memory, and Ubuntu 22.04.4 LTS.
@@
Smooth loss sets top $N$ at 5. Focal loss contains $\alpha$ at 0.6 and $\gamma$ at 2.0. Dynamic loss has $N$ at 5, $T_{D}$ at 0.1, and $\alpha$ at 5. Additive loss has $\alpha$ at 0.7, and Common loss has $\beta$ at 0.2. Self Consistency decoding strategy sets $N_{\text{SCD}}$ at 3, and $\mathcal{D}_{\text{Optimal}}$ at 0.7. The seed is set at 42, and we utilize fp16 for faster training.